A short run cost function for an entrepreneur is

q3 − 8q2 + 30q + 60 Determine the

price at which the entrepreneur cases production in an ideal market. Also, derive the

supply function.


1
Expert's answer
2021-03-02T04:50:08-0500
"MC=3q^2-16q+30"

"MR=MC=>p=3q^2-16q+30"

"p=3(q^2-2(\dfrac{8}{3})q+\dfrac{64}{9})-\dfrac{64}{3}+30"

"\dfrac{p}{3}=(q-\dfrac{8}{3})^2+\dfrac{26}{9}"

"(q-\dfrac{8}{3})^2=\dfrac{p}{3}-\dfrac{26}{9}"

"q-\dfrac{8}{3}=\sqrt{\dfrac{p}{3}-\dfrac{26}{9}}"

Supply function

"q=\sqrt{\dfrac{p}{3}-\dfrac{26}{9}}+\dfrac{8}{3}"